if the scale ratio of creating a house on a blueprint is 1 1/2 inch = 10 feet how many feet is it if you have 2 inches on the blue print

Answer by checkley79(3341)

1.5/10=2/X
1.5X=2*10
1.5X=20
X=20/1.5
=13.33 FEET ANS.